Grade 12th passGeneral Physics

A river 2Km wide is flowing at the rate of 2km/hr.A boatman who can row a boat at the speed of 4km/hr in still water goes a distance of 2Km upstream and then comesback. The time taken by him to complete his journey.

Parul Rajput
9 Years agoGrade 12th pass

1 Answer

deepak
9 Years ago
let the velocity of the water be 2\widehat{i}
and the boats velocity be a\widehat{i}+b\widehat{j} = 4km/hr
 
for the boat to cross in a straight line the \widehat{i} component of its velocity must cancel out the velociity of the water
 
hence a=-2
 
as we know
 
4= \sqrt{a^2 + b^2}
therefore b= 2\sqrt{3}
 
hence the time taken to cover a distance of 4km in the y direction = 2 / \sqrt{3} sec
 
hope this was useful